# Supp_Appendix_v4

### v3

@@ 4.2   Using Internal Validity to Bound Average Effects
Substituting equation (4.5) in (4.6) and assuming that $\mathbb{E}[Y(1)-Y(0)\mid W^{-}=1]$ and $\mathbb{E}[Y(1)-Y(0)\mid W^{+}+W^{-}=0]$ lie in $[B_{\ell},B_{u}]$ yields
@@ 4.3   Quantifying Internal Validity Given $\tau_{0}$
To illustrate this theorem, let $\tau_{0}(X)$ be continuously distributed with support $[\underline{\tau},\overline{\tau}]$, and suppose $\mu_{0}\in[\underline{\tau},\overline{\tau}]$. Without loss of generality, assume $E_{0}\geq\mu_{0}$. If $E_{0}=\mu_{0}$, then the estimand is perfectly representative of the population since it equals the average treatment effect over it. In the case where $E_{0}>\mu_{0}$, the estimand is not representative of the entire population. We are searching for the largest subpopulation $\{W^{*}=1\}$ such that $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu_{0}$. Initializing $W^{*}$ at $W_{0}$, removing the subpopulation with the largest values of $\tau_{0}(x)$ yields the steepest decrease in $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$. Therefore, $\overline{P}(a,W_{0};\{\tau_{0}\})$ is obtained by removing a subpopulation of the kind $W^{-}(\alpha)=\mathbbm{1}(\tau_{0}(X)>\alpha)\cdot W_{0}$ for a given threshold $\alpha$. This threshold is determined by the constraint
@@ 5.1   Unconfoundedness
Suppose the target estimand is the average treatment effect, i.e., $W_{0}=1$ almost surely. By Theorem 3.1, there exists a regular subpopulation $W^{*}$ such that $\beta_{\text{OLS}}$ equals the average treatment effect over $W^{*}$ since the weight function $a(X)=p(X)(1-p(X))$ is nonnegative. By Theorem 4.1, the upper bound on the size of subpopulation $W^{*}$ is
@@ 5.3   Difference-in-Differences
We now consider the weights obtained in Proposition 5.4 under its assumptions. These weights are nonnegative and therefore Theorem 3.1 guarantees the existence of a causal representation for $\beta_{\text{TWFE}}$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$. Using Theorem 4.1, the internal validity of $\beta_{\text{TWFE}}$ relative to target parameter $\mathbb{E}[Y(1)-Y(0)\mid D=1]$ is given by
@@ 6   Estimation and Inference
Here $c_{n}$ is a tuning parameter that converges to 0 as $n$ diverges. We start by noting that we can estimate $\mathbb{E}[a(X)\mid W_{0}=1]$ via $\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})}$, which will be consistent under standard conditions on $\widehat{a}(\cdot)$ and $\widehat{w}_{0}(\cdot)$. Estimating $a_{\max}=\sup(\text{supp}(a(X)\mid W_{0}=1))$ is more delicate. In some of our examples, this supremum is known or can be bounded above without using data. For example, the OLS estimand under unconfoundedness has weights $a(X)=\operatorname{var}(D\mid X)$ which are bounded above by $1/4$. If $X$ is continuously distributed, then 1/2 may lie in the support of $p(X)$, and thus we may avoid the estimation of $a_{\max}$. The IV estimand of Section 5 has weights $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)$ which are similarly bounded above by $1/4$. Similarly, $a_{\text{2SLS}}(X)=|\operatorname{cov}(D,Z\mid X)|\leq 1/4$ by the Cauchy–Schwarz inequality. If knowledge of $a_{\max}$ is not assumed, but $\text{supp}(X\mid W_{0}=1)$ is known and $a(x)$ is continuous, [^4] Note that $a(x)$ is trivially continuous on finite support. then $\sup_{x\in\text{supp}(X\mid W_{0}=1)}\widehat{a}(x)$ will be consistent for $a_{\max}$ when $\widehat{a}(x)$ is consistent for $a(x)$ uniformly in $x\in\text{supp}(X\mid W_{0}=1)$. Many parametric and nonparametric estimators for $a(\cdot)$ satisfy this requirement.
@@ 7   Empirical Application
The conclusion changes, however, when we explicitly target the average treatment effect on the treated (ATT), that is, the average effect for the largest subpopulation for which such an effect is identified under standard assumptions. Using the approach of Wooldridge (2021), we obtain an estimate of –5.530 with a $p$-value of 0.138. The approach of Callaway and Sant’Anna (2021) produces an estimate of –10.220 and a $p$-value of 0.001. These estimates are more strongly suggestive of a causal effect of unilateral divorce laws than the TWFE estimate.
{adjustwidth}
-1.25in-1.25in A. Estimates of the effects of unilateral divorce laws TWFE ATT Callaway and Sant’Anna Wooldridge –0.604 –10.220 –5.530 (2.622) (3.086) (3.650) B. Internal validity of the TWFE estimand based on Proposition 5.3 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0 0.3873 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0 0.6216 C. Internal validity of the TWFE estimand based on Proposition 5.4 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0.1400 0.4802 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0.2246 0.7707
@@
While the TWFE estimate and the two estimates of the ATT are quite different, this paper focuses on another implication of the nonuniformity of the TWFE weight function. We ask: How representative of the underlying population is the TWFE estimand? What is the internal validity of this estimand if we are interested in the treated subpopulation? Panel B of Table 1 reports our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)$, based on the representation of the TWFE estimand in de Chaisemartin and D’Haultfœuille (2020), revisited in our Proposition 5.3. First, because the weights on some group-time average treatment effects are negative, the TWFE estimand does not have a causal interpretation uniformly in $\tau_{0}$, ^$\widehat{\mathbb{P}}(W^{*}=1)=\widehat{\mathbb{P}}(W^{*}=1\mid D=1)=0$. Second, when we estimate the CATE function and use these estimates in constructing the bounds, we conclude that the TWFE estimand corresponds to the average treatment effect for at most 62.16% of the treated units or 38.73% of the entire population.
@@ Appendix A Proofs of Key Results
()($\Longrightarrow$)( ⟹ ) First, suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$. By Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{all}}$. Let $\tau^{*}(X)=\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{P}(W^{*}=1\mid W_{0}=1)}$. We have $\mathbb{E}[a(X)^{2}]<\infty$ by Assumption 3.1 and $\mathbb{E}[\underline{w}^{*}(X)^{2}]\leq 1$, which holds by construction. Hence, $\mathbb{E}[\tau^{*}(X)^{2}]<\infty$, and thus $\tau^{*}\in\mathcal{T}_{\text{all}}$.
@@
11 11 $=\mathbb{P}(\tau^{*}(X)=0\mid W_{0}=1)=\mathbb{P}\left(a(X)=\frac{\underline{w}^{*}(X)\mathbb{E}[a(X)\mid W_{0}=1]}{\mathbb{P}(W^{*}=1\mid W_{0}=1)}\mid W_{0}=1\right).$ (A.2)
@@
()($\Longrightarrow$)( ⟹ ) First, suppose $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$ and suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$. Since $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$, we can without loss of generality suppose that $\mathbb{P}(\tau_{0}(X)>\mu_{0}\mid W_{0}=1)=1$. Since $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$, we can write by Proposition 3.1
@@
()($\Longrightarrow$)( ⟹ ) First, we suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$. By Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{lin}}$. Therefore,
00 0 $=\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})$
@@
()($\Longrightarrow$)( ⟹ ) First, let $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$. We will show that $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BD}}(K))=\emptyset$ by way of contradiction.
@@ Appendix B Weakly Causal Estimands and Uniform Causal Representations in $\mathcal{T}_{\text{all}}$
Thus, an estimand is weakly causal if all CATEs having the same sign implies the estimand also has that sign. Whether an estimand satisfies this condition also depends on $\mathcal{M}$, the set of allowed ASFs. To compare weak causality to our result on uniform causal representations, we consider $\mathcal{M}_{\text{all}}\coloneqq\mathbb{R}^{2K}$, the unrestricted set of ASFs. The corresponding unrestricted set of CATE functions, which we denoted by $\mathcal{T}_{\text{all}}$, allows $\tau$ to be any vector in $\mathbb{R}^{K}$. We consider estimands characterized by our equation (1.1). We note that these estimands rule out “level dependence,” i.e., that the estimand changes if potential outcomes $(Y(0),Y(1))$ are translated to $(Y(0)+c,Y(1)+c)$ for some constant $c\in\mathbb{R}$. For example, the IV estimand is generally level dependent when the propensity score $\mathbb{P}(Z=1\mid X)$ is nonlinear in $X$ [^6] See p. 17 in BBMT. With these choices, we can show the two definitions are equivalent.
@@ Appendix C Details on Estimation and Inference
Let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, denote by $p_{j}=\mathbb{P}(X=x_{j})$ the frequency of cell $j$, and let $\widehat{p}_{j}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})$ denote its sample frequency. Let $\widehat{\theta}=(\widehat{\textbf{a}},\widehat{\textbf{w}}_{0},\widehat{\textbf{p}})$ where $\widehat{\textbf{a}}=(\widehat{a}(x_{1}),\ldots,\widehat{a}(x_{K}))$, $\widehat{\textbf{w}}_{0}=(\widehat{w}_{0}(x_{1}),\ldots,\widehat{w}_{0}(x_{K}))$, and $\widehat{\textbf{p}}=(\widehat{p}_{1},\ldots,\widehat{p}_{K})$. Let $\theta=(\textbf{a},\textbf{w}_{0},\textbf{p})$ denote their population counterparts.
@@
[…]
@@
[…]
@@
[…]
@@
as $n\rightarrow\infty$, where $\widehat{\psi}$ is defined by
$\widehat{\psi}(\mathbb{Z}^{*})\;$ $=\;\sum_{j=1}^{K}\frac{\widehat{w}_{0}(x_{j})\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\widehat{a}_{\max}}\mathbb{Z}^{*}_{\textbf{a}}(j)\;-\;\frac{\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]}{\widehat{a}_{\max}^{2}}\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}\mathbb{Z}^{*}_{\textbf{a}}(j)$
@@
where $\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})}$, ^$\widehat{\mathbb{P}}(W_{0}=1)=\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})$, $\widehat{a}_{\max}=\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})$, and
@@
    For bootstrap samples $b=1,\ldots,B$, compute $\widehat{\theta}^{\ast,b}=(\widehat{\textbf{a}}^{\ast,b},\widehat{\textbf{w}}_{0}^{\ast,b},\widehat{\textbf{p}}^{\ast,b})$ and $\mathbb{Z}^{\ast,b}=\sqrt{n}(\widehat{\theta}^{\ast,b}-\widehat{\theta})$;
@@
Computing a one-sided confidence interval for $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ of the kind $[0,\widehat{\overline{P}}^{+}]$ can be cast as doing inference on intersection bounds. Chernozhukov, Lee, and Rosen (2013) offer methods for such problems. Equivalently, the computation of a one-sided confidence interval $[\widehat{\overline{P}}^{-},1]$ is related to inferential questions in union bounds: see Bei (2024). We leave all details for future work.
@@
By the definition of $\alpha^{+}$ as an infimum, we must have that $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, implying that $R(\alpha)$ is discontinuous at $\alpha^{+}$. By Lemma D.1, $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]$ is left-continuous and satisfies $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]\leq R(\alpha)$. Therefore, since $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, we must have that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+}-\varepsilon)\mid W_{0}=1]<0$ for all $\varepsilon>0$. Letting $\varepsilon\searrow 0$ yields that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]\leq 0$. Therefore $-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]/(\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1))\geq 0$ and $\underline{w}^{+}(X)\geq 0$.
@@
[…]

### v4

@@ 4.2  Using Internal Validity to Bound Average Effects
Substituting equation (4.5) in (4.6) and assuming that $\mathbb{E}[Y(1)-Y(0)\mid W^{-}=1]$ and $\mathbb{E}[Y(1)-Y(0)\mid W^{+}+W^{-}=0]$ lie in $B_{\ell},B_{u}]$ yields
@@ 4.3  Quantifying Internal Validity Given $\tau_{0}$
To illustrate this theorem, let $\tau_{0}(X)$ be continuously distributed with support $\underline{\tau},\overline{\tau}]$, and suppose $\mu_{0}\in[\underline{\tau},\overline{\tau}]$. Without loss of generality, assume $E_{0}\geq\mu_{0}$. If $E_{0}=\mu_{0}$, then the estimand is perfectly representative of the population since it equals the average treatment effect over it. In the case where $E_{0}>\mu_{0}$, the estimand is not representative of the entire population. We are searching for the largest subpopulation $\{W^{*}=1\}$ such that $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu_{0}$. Initializing $W^{*}$ at $W_{0}$, removing the subpopulation with the largest values of $\tau_{0}(x)$ yields the steepest decrease in $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$. Therefore, $\overline{P}(a,W_{0};\{\tau_{0}\})$ is obtained by removing a subpopulation of the kind $W^{-}(\alpha)=\mathbbm{1}(\tau_{0}(X)>\alpha)\cdot W_{0}$ for a given threshold $\alpha$. This threshold is determined by the constraint
@@
Figure 2: Characterizing a Representative Subpopulation When $\tau_{0}$ Is Known
Note: The figure assumes that $\tau_{0}(X)$ is continuously distributed, that $W_{0}=1$ almost surely, and that $\mu<\mathbb{E}[\tau_{0}(X)]=\text{ATE}$.
Theorem 4.2 can also be illustrated visually. In Figure 2, the probability density function of $\tau_{0}(X)$ is drawn. In this figure, it is assumed that $\tau_{0}(X)$ is continuously distributed, that $W_{0}=1$ almost surely, and that $\mu<\mathbb{E}[\tau_{0}(X)]=\text{ATE}$. The representative subpopulation is obtained by trimming away covariate values that correspond to $\tau_{0}(X)\geq\alpha^{+}$, where $\alpha^{+}$ is determined by the equation $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq\alpha^{+}]=\mu$. The size of the shaded area is the measure of internal validity.
@@ 5.1  Unconfoundedness
Suppose the target estimand is the average treatment effect, i.e., $W_{0}=1$ almost surely. By Theorem 3.1, there exists a regular subpopulation $W^{*}$ such that $\beta_{\text{OLS}}$ equals the average treatment effect over $W^{*}$ since the weight function a(X)=p(X)(1p(X))a(X)=p(X)(1-p(X)) is nonnegative. By Theorem 4.1, the upper bound on the size of subpopulation $W^{*}$ is
@@ 5.3  Difference-in-Differences
We now consider the weights obtained in Proposition 5.4 under its assumptions. These weights are nonnegative and therefore Theorem 3.1 guarantees the existence of a causal representation for $\beta_{\text{TWFE}}$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$. [^5] In the context of Proposition 5.4, $\tau_{0}$ is a function of $G$ only, thus $\mathcal{T}_{\text{all}}$ denotes the set of all functions of $G$ with finite second moments. Note that this is a strict subset of all “time-heterogeneous” conditional average treatment effects, $\mathbb{E}[Y(1)-Y(0)\mid D=1,G,P]$. Using Theorem 4.1, the internal validity of $\beta_{\text{TWFE}}$ relative to target parameter $\mathbb{E}[Y(1)-Y(0)\mid D=1]$ is given by
@@ 6  Estimation and Inference
[…]
@@ 7  Empirical Application
The conclusion changes, however, when we explicitly target the average treatment effect on the treated (ATT), that is, the average effect for the largest subpopulation for which such an effect is identified under standard assumptions. Using the approach of Callaway and Sant’Anna (2021), we obtain an estimate of –10.220 with a $p$-value of 0.001. The approach of Wooldridge (2025) produces an estimate of –5.530 and a $p$-value of 0.138. These estimates are more strongly suggestive of a causal effect of unilateral divorce laws than the TWFE estimate.
@@
A. Estimates of the effects of unilateral divorce laws
TWFE ATT
Callaway and Sant’Anna Wooldridge
–0.604 –10.220 –5.530
(2.622) (3.086) (3.650)
B. Internal validity of the TWFE estimand based on Proposition 5.3
uniformly in $\tau_{0}$ given $\tau_{0}$
$\widehat{\mathbb{P}}(W^{*}=1)$ 0 0.3873
$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0 0.6216
C. Internal validity of the TWFE estimand based on Proposition 5.4
uniformly in $\tau_{0}$ given $\tau_{0}$
$\widehat{\mathbb{P}}(W^{*}=1)$ 0.1400 0.4802
$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0.2246 0.7707
@@ A. Estimates of the effects of unilateral divorce laws
While the TWFE estimate and the two estimates of the ATT are quite different, this paper focuses on another implication of the nonuniformity of the TWFE weight function. We ask: How representative of the underlying population is the TWFE estimand? What is the internal validity of this estimand if we are interested in the treated subpopulation? Panel B of Table 1 reports our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)$, based on the representation of the TWFE estimand in de Chaisemartin and D’Haultfœuille (2020), revisited in our Proposition 5.3. First, because the weights on some group-time average treatment effects are negative, the TWFE estimand does not have a causal interpretation uniformly in $\tau_{0}$, $\widehat{\mathbb{P}}(W^{*}=1)=\widehat{\mathbb{P}}(W^{*}=1\mid D=1)=0$. Second, when we estimate the CATE function and use these estimates in constructing the bounds, we conclude that the TWFE estimand corresponds to the average treatment effect for at most 62.16% of the treated units or 38.73% of the entire population.
@@ Appendix A Proofs of Key Results
$(\Longrightarrow)$ First, suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$. By Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{all}}$. Let $\tau^{*}(X)=\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{P}(W^{*}=1\mid W_{0}=1)}$. We have $\mathbb{E}[a(X)^{2}]<\infty$ by Assumption 3.1 and $\mathbb{E}[\underline{w}^{*}(X)^{2}]\leq 1$, which holds by construction. Hence, $\mathbb{E}[\tau^{*}(X)^{2}]<\infty$, and thus $\tau^{*}\in\mathcal{T}_{\text{all}}$.
@@
$1$ $=\mathbb{P}(\tau^{*}(X)=0\mid W_{0}=1)=\mathbb{P}\left(a(X)=\frac{\underline{w}^{*}(X)\mathbb{E}[a(X)\mid W_{0}=1]}{\mathbb{P}(W^{*}=1\mid W_{0}=1)}\mid W_{0}=1\right).$ (A.2)
@@
$(\Longrightarrow)$ First, suppose $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$ and suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$. Since $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$, we can without loss of generality suppose that $\mathbb{P}(\tau_{0}(X)>\mu_{0}\mid W_{0}=1)=1$. Since $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$, we can write by Proposition 3.1
@@
$(\Longrightarrow)$ First, we suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$. By Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{lin}}$. Therefore,
$0$ $=\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})$
@@
$(\Longrightarrow)$ First, let $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$. We will show that $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BD}}(K))=\emptyset$ by way of contradiction.
@@ Appendix B Weakly Causal Estimands and Uniform Causal Representations in $\mathcal{T}_{\text{all}}$
Thus, an estimand is weakly causal if all CATEs having the same sign implies the estimand also has that sign. Whether an estimand satisfies this condition also depends on $\mathcal{M}$, the set of allowed ASFs. To compare weak causality to our result on uniform causal representations, we consider $\mathcal{M}_{\text{all}}\coloneqq\mathbb{R}^{2K}$, the unrestricted set of ASFs. The corresponding unrestricted set of CATE functions, which we denoted by $\mathcal{T}_{\text{all}}$, allows $\tau$ to be any vector in $\mathbb{R}^{K}$. We consider estimands characterized by our equation (1.1). We note that these estimands rule out “level dependence,” i.e., that the estimand changes if potential outcomes $(Y(0),Y(1))$ are translated to $(Y(0)+c,Y(1)+c)$ for some constant $c\in\mathbb{R}$. For example, the IV estimand is generally level dependent when the propensity score $\mathbb{P}(Z=1\mid X)$ is nonlinear in XX. [^8] See p. 17 in BBMT. With these choices, we can show the two definitions are equivalent.
@@ Appendix C Details on Estimation and Inference
[…]
@@
$\sqrt{n}(\widehat{\theta}-\theta)\xrightarrow{d}\mathbb{Z}$
@@
$\sqrt{n}(\widehat{\overline{P}}-\overline{P})$ $\xrightarrow{d}\psi(\mathbb{Z}),$
@@
Let $\mathbb{Z}^{*}\coloneqq(\mathbb{Z}^{*}_{\textbf{a}},\mathbb{Z}^{*}_{\textbf{w}_{0}},\mathbb{Z}^{*}_{X})\in\mathbb{R}^{3K}$ be a random vector such that $\mathbb{Z}^{*}\overset{p}{\rightsquigarrow}\mathbb{Z}$, where $\overset{p}{\rightsquigarrow}$ denotes convergence in probability conditioning on the data used to compute $\widehat{\theta}$.
@@
$\widehat{\psi}(\mathbb{Z}^{*})\overset{p}{\rightsquigarrow}\psi(\mathbb{Z})$
as $n\rightarrow\infty$, where $\widehat{\psi}$ is defined by
@@
where $\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})}$, $\widehat{\mathbb{P}}(W_{0}=1)=\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})$, $\widehat{a}_{\max}=\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})$, and
@@
[…]
@@
Computing a one-sided confidence interval for $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ of the kind $0,\widehat{\overline{P}}^{+}]$ can be cast as doing inference on intersection bounds. Chernozhukov, Lee, and Rosen (2013) offer methods for such problems. Equivalently, the computation of a one-sided confidence interval $\widehat{\overline{P}}^{-},1]$ is related to inferential questions in union bounds: see Bei (2024). We leave all details for future work.
@@
By the definition of $\alpha^{+}$ as an infimum, we must have that $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, implying that $R(\alpha)$ is discontinuous at $\alpha^{+}$. By Lemma D.1, $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]$ is left-continuous and satisfies $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]\leq R(\alpha)$. Therefore, since $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, we must have that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+}-\varepsilon)\mid W_{0}=1]<0$ for all $\varepsilon>0$. Letting $\varepsilon\searrow 0$ yields that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]\leq 0$. Therefore $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]/(\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1))\geq 0$ and $\underline{w}^{+}(X)\geq 0$.
@@
By construction, $W^{+}\in\{0,1\}$, $\mathbb{P}(W^{+}=\;1\mid X,W_{0}\;=\;1)\;=\;\underline{w}^{+}(X)$, $\mathbb{P}(W_{0}\;=\;1\mid W^{+}=\;1)\;=\;1$, and $W^{+}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$. Also, since $\mu_{0}\in\mathcal{S}(\tau_{0};W_{0})$, $\mathbb{P}(T_{\mu}\leq 0\mid W_{0}=1)=\mathbb{P}(\tau_{0}(X)\leq\mu_{0}\mid W_{0}=1)>0$. Since $\alpha^{+}>0$ we have that $\mathbb{P}(W^{+}=1\mid W_{0}=1)\geq\mathbb{P}(T_{\mu}<\alpha^{+}\mid W_{0}=1)\geq\mathbb{P}(T_{\mu}\leq 0\mid W_{0}=1)>0$. Therefore $W^{+}$ is a regular subpopulation of $W_{0}$ for which $\mathbb{P}(W^{+}=1\mid W_{0}=1)=P^{+}$, hence $P^{+}$ is the maximum.